OpenStudy (anonymous):

Please help me! 4x/x+3-4=x-8/x-3. How do I find the answers to this? There should be two, I just can't figure out how to solve this with the stupid -4 :(

OpenStudy (anonymous):

It says that isnt right

OpenStudy (anonymous):

oh crud I messed up lol sorry

OpenStudy (anonymous):

4x/(x + 3 - 4) = (x-8)/(x-3) = 4x(x-3)(x + 3 - 4)/(x + 3 - 4) = (x + 3 - 4)(x-8)(x-3)/(x-3) = 4x(x-3) = (x + 3 - 4))(x-8)

OpenStudy (anonymous):

the answer should look like this: number, number

OpenStudy (anonymous):

-4 is seperate

OpenStudy (anonymous):

elimante the fractions on both sides by mutliplying both sides of the equation by the denominators

OpenStudy (anonymous):

4x(x-3) = (x + 3 - 4)(x-8) 4x^(2) - 12 = -8x + -24 + 32 + x^(2) + 3x - 4x

OpenStudy (anonymous):

4x^(2) - 12 = -8x + -24 + 32 + x^(2) + 3x - 4x = 4x^(2) - 12 = -9x + 8 + x^(2) = 0 = -9x + 8 + x^(2) - 4x^(2) + 12 0 = -3x^(2) - 9x + 20

OpenStudy (anonymous):

now factor

OpenStudy (anonymous):

oh crud I didn't look at your problem lol

OpenStudy (anonymous):

(4x/(x+3)) + 4(x+3)/(x+3) = (4x + 4(x+3))/(x+3)

OpenStudy (anonymous):

now just do what I did before to solve

OpenStudy (anonymous):

you can multiply the top and the botom of a fraction by the same number (excluding 0) and it wont change the value of the fraction this can be used to add whole numbers to fractions. To prove this 4(x-1)/(x-1) = 4

OpenStudy (anonymous):

Because (x-1)/(x-1) = 1 and 4*1 = 4

OpenStudy (anonymous):

(4x + 4(x+3))/(x+3) = x-8/x-3 Can you solve this now?

OpenStudy (anonymous):

yes no maybe?

OpenStudy (anonymous):

(4x + 4(x+3))/(x+3) = (x-8)/(x-3) = (4x + 4)(x-3) = (x-8)(x+3)

OpenStudy (anonymous):

do the substraction for the LHS then do opposite multiplication from the eq given by @cuddlepony : (4x-4(x+3)) (x-3) = (x+3)(x-8)

OpenStudy (anonymous):

I got that wrong too, it was -12,5 :(

OpenStudy (anonymous):

show your work

OpenStudy (anonymous):

and we will show you where you went wrong